# THOMAS et al

### v1

@@ Abstract
[Summary]We make use of Kronecker structure for scaling Gaussian Process models to large-scale, heterogeneous, clinical data sets. Repeated measures, commonly performed in clinical research, facilitate computational acceleration for nonlinear Bayesian nonparametric models and enable exact sampling for non-conjugate inference, when combinations of continuous and discrete endpoints are observed. Model inference is performed in Stan, and comparisons are made with brms on simulated data and two real clinical data sets, following a radiological image quality theme. Scalable Gaussian Process models compare favourably with parametric models on real data sets with 17,460 observations. Different GP model specifications are explored, with components analogous to random effects, and their theoretical properties are described.
[…]
@@
articletype: Article Type00footnotetext: Abbreviations: GP, Gaussian Process; HMC, Hamiltonian Monte Carlo; RCT, Randomised Control Trial; ARD, Automatic Relevance Determination; NUTS, No U-Turn Sampler; CTA, Computed Tomography Angiography; keV, kiloelectronVolts; ROI, Region Of Interest; SVM, Support Vector Machine; ESS, Effective Sample Size; HU, Hounsfield Unit; ICM, Intrinsic Coregionalization Model
@@ 1 Introduction
Here we present and run new GP models using repeated measurements in the covariate structure and multiple heterogeneous (mixed continuous and discrete) outputs: exact inference in Stanstancore is scaled to tens of thousands of measurements on the processors of a domestic-issue laptop. The models represent the joint covariance between all of the data as Kronecker-structured, and perform Hamiltonian Monte Carlo (HMC) sampling for hyperparameters, missing output values, and on the latent GP space for non-conjugate inference in the presence of heterogeneous outputs. The clinical data sets follow the theme of radiological image quality, in which the covariates consist of patient characteristics, body locations and time, while the outputs represent measurements of image quality, either continuously "objectively" on the Hounsfield scalehounsfieldunits of radiodensity, or discretely "subjectively" from expert evaluations.
{boxtext}
@@
  • An implementation of the relevant Gaussian Process models in Stanstancore, rstan (with wrapper in R), allowing for running the inference on multiple real world data sets, and benchmarking various model specificatons such as covariance function choices.
@@ 2.2 Gaussian Processes
$y|x=$ $f(x)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (2)
Specific choices of covariance functions $k(x,x^{\prime},\theta)$ have corresponding implicit parametric basis functions. One advantage of the "function-space" formulation is the ability to use a finite representation of a function with a potentially infinite-dimensional parametric representation: this is the sense in which the models are considered "nonparametric", in that they avoid specifying a parametric model for the latent mean function.
@@ 2.3 Scalability and Kronecker-Structure Covariance Matrices
Asserting a GP with a covariance function $k(x,x^{\prime},\theta)$ over a data set with $N$ covariate observations defines a $N\times N$ covariance222also known as kernel matrix matrix $\mathbf{K}$, where the matrix element $\mathbf{K}[i,j]$ is equal to the covariance function $k(x^{i},{x^{j}}^{\prime},\theta)$ evaluated at the $ith$ and $jth$ data point. One challenge of working with GPs is the need to perform a decomposition of the $N\times N$ covariance matrix, which is the dominant computational demand when evaluating the marginal likelihood for sampling or computing the predictive distributions. This results in $\mathcal{O}(N^{3})$ cubic scaling in computational costs with the number of data points $N$, ruling out exact inference for generic covariance matrices for larger data sets . Various methods exist to enable approximate inference for larger data setsliu2020gaussian, thomas2017scalable, while exact inference is possible for larger data sets when there is structure in the covariance matrix that can be exploited, such as Kronecker or Toeplitzsaatcci2012scalable. In instances where data can be represented as lying in a grid structure, with separable covariance structure between each dimension of the grid, the full covariance matrix can be represented as a Kronecker product between the dimensions. A Kronecker product is an operation performed on two matrices generating a third matrix composed of each individual element of the first matrix separately multiplied with the entire second matrix, combined in a blockwise fashion, i.e. for two matrices $A$ and $B$:
@@
For a data set consisting of covariates measured over a two dimensional grid, for example the pixels of an image, the covariance matrix over the entire data set ($\mathbf{K}_{12}=K([x_{1},x_{2}],[x^{\prime}_{1},x^{\prime}_{2}])$) can be represented as the Kronecker product between two covariance matrices representing each grid dimension independently ($\mathbf{K}_{1}=K(x_{1}.x^{\prime}_{1})$, $\mathbf{K}_{2}=K(x_{2}.x^{\prime}_{2}),$), with the constraint that the covariance functions used are separable, i.e. for a grid composed of $n_{1}$ vertical grid points $x_{1}$, and $n_{2}$ horizontal grid points $x_{2}$, then:
@@
We can apply this reasoning to clinical scenario described in 2.1: for a longitudinal study consisting of $N_{1}$ individuals measured at $N_{2}$ time points at $N_{3}$ anatomical locations, the covariate data X can be represented as lying on a three dimensional grid, where one of the grid dimensions $x_{1}$ is the variation between individuals, the second $x_{2}$ as the variation in time, and the third $x_{3}$ variation in anatomical location. Consequently, defining separate covariance matrices representing the variation across individuals (the $N_{1}\times N_{1}$ matrix $\mathbf{K}_{1}$), variation over time (the $N_{2}\times N_{1}$ matrix $\mathbf{K}_{2}$), and anatomical variation (the $N_{3}\times N_{3}$ matrix $\mathbf{K}_{3}$) we can represent the covariance between the covariate measurements $X$ as:
@@ 2.4 Heterogeneous Multi-Task Gaussian Processes
The GP framework extends naturally to multi-dimensional responses $\mathbf{Y}$, analogous to the case of multivariate regressionalvarez2012kernels, bonilla2007multi. The different response dimensions (also known as "tasks" or "outputs") are appended into a single vector, and a between-task covariance matrix $\mathbf{K}_{out}$ is used to model correlations between the tasks. When the different tasks are evaluated at the same values of the covariates, generating a covariance matrix between covariates $\mathbf{K}_{X}$ then Kronecker structure again emerges in the full covariance matrix $\mathbf{K}_{f}=\mathbf{K}_{out}\otimes\mathbf{K}_{X}$, and computational accelerations become possible.
@@
All of the data sets used in this article use one Kronecker component for the multi-output correlation, and three Kronecker grid components for the covariates, which are grouped into grid dimensions $x_{1}$, $x_{2}$ and $x_{3}$. The full covariance structure is therefore:
@@ 2.5 Covariance Matrix Design and Random Effects
Random effects are often desirable in the presence of repeated measurements, where we are not necessarily interested in the variation between individuals but we would still like to include it in the model gelman2007data. This can be performed in an elegant way within GP regression models by including a structured diagonal noise component representing the random variation across individuals. In the Kronecker context, this can achieved by adding spherical noise to the Kronecker product component that represents the variation between individuals, representing the "noise" sampled when moving between individuals. As this is an additive covariance function where one of the covariance functions is diagonal noise, the same problem emerges when combining additive covariance functions with Kronecker structure: the resulting covariance function components cannot be interpreted totally straightforwardly, as they are multiplied with the covariance functions associated with the other Kronecker components. Formally, a traditional random effects model is represented in the first half of the following inequality, and the model we implemented in the second:
@@
This formulation of a random effect is equivalent to adding some extra variance to the coefficients corresponding to non-patient specific covariates. We detail this in Appendix B, for the case where all kernels are linear and thus GP regression is equivalent to Bayesian linear regression.
We ran separate models with and without the "random effect" noise in the individual-level covariance function to explore its influence on the model fit and predictions. We call the models with and without this the mixed-effect GP ("GP.m") and the fixed-effect GP ("GP.f").
@@ 2.7 Stan Implementation
The methods described here were implemented in Stanstancore, a probabilistic programming language designed for Bayesian inference, in which model specification is performed explicitly within the language, and No U-Turn Sampler (NUTS) Hamiltonian Monte Carlohoffman2014no (HMC) is performed for the model parameters. The Gaussian Process was represented for inference with a standardised, uncorrelated Gaussian latent variable $\eta$, which was reshaped and transformed by the covariance Kronecker components to form the latent variable $f$ conditioned on the covariance. HMC was performed for the GP hyperparameters and noise terms, the latent function representation $\eta$, the latent function $f$ for discrete-valued tasks, and the missing values of the outputs. The Stan programs were called from within the R programming language via Rstanrstan. Similar computational speedups have been achieved in Stan previously flaxman2015fast, Ronneberg2021.
@@ 2.8 Missing Data
The inference for continuous-valued missing outputs $y$ can be integrated in a straightforward way with most inference schemes. Given that Stan cannot perform inference over discrete variables, then the missing binary outputs pose a challenge. This was resolved by calculating the log-likelihood contributions according to the Bernoulli likelihood with an appropriate link function, with the missing $y$ varying continuously between zero and one. This allows for efficient gradient-informed inference procedures, and the interpretation of the imputed variables as probabilities.
@@ 2.9 Radiological Image Quality
The clinical theme of this article is radiological image quality, a field in which it is common to combine continuous-valued "objective" measurements of image quality with discrete-valued "subjective" expert evaluations of image quality. This article uses two real data sets, the first of which is observational time-series data, and the second of which comes from a RCT with thorough repeated measures. In each case the effect of primary interest is the effect on the image quality of volume of contrast medium per body weight.
The first real data set was previously published as a clinical study in tran2022indirect. The image quality was measured in venal blood vessels of 53 patients, with continuous-valued attenuation in Hounsfield Units as the objective endpoint, and binary evaluations of image quality from three different consultant radiologists as the subjective endpoints. Each patient was injected with a different quantity of contrast medium per body weight, and evaluations of the images were taken at six thirty-second interval time points after injection. Evaluations were also performed for each side of the body. Further covariates used were gender, age and tube voltage. Further clinical details can be found here. The combination of 53 patients, 4 endpoints, 7 time points and 2 body sides implies a covariance structure of size 2,968, but the use of the Kronecker product for each of these contributions means that the computations are readily tractable on a personal laptop.
The second real data set comes from a Randomised Control Trial assessing image quality in arterial blood vessels following Computed Tomography Angiography (CTA) kristiansen2023halved. 210 patients were randomised to receive full versus half doses of contrast volume per body weight, with measurements repeated at 7 different locations in the body, and 3 different spectral energy levels measured in kiloelectronVolts (keV). Examples of images recorded at different body levels and energy levels are presented in Figure 1. Attenuation level and image noise recorded in selected Regions Of Interest (ROIs), each measured in Hounsfield units were used as continuous-valued "objective" outcomes, and two consultant radiologists evaluated every image on a nominal scale for image quality. Because of the very skew distribution of the nominal data, these were simplified to a binary variable of "Excellent" vs every other level, results into two binary "subjective" endpoints. Examples of the ordinal scale of subjective image quality are shown in Figure 2. Sex, age and flow speed of injection were also collected as relevant covariates. The combination of 210 patients, 4 endpoints, 7 body locations and three energy levels resulted in a data set of size $N=17,640$ but again the Kronecker structure enabled exact inference with limited computational resources.
@@ 3 Related Work
Gaussian Processes have been used for decades under various names such as kriging or Bayesian kernel regression, and are related to the widely used Support Vector Machines (SVMs). Their used has increased in the past couple of decades with the advent of greater computational resources, and much research has occurred concerning their scalability to larger data sets under contemporary computational constraints. These methods often include variational inference schemes via the "inducing point" frameworkhensman2013gaussian, or spectral methods to approximate the full covariance function via sampled Fourier featuresrahimi2007random. The speedups possible through Kronecker structure have been used for some time, but have previously been used on structured image datagilboa2014image, drug combinationsRonneberg2021, Ronneberg2023, spatio-temporal modellingGelfand2004, and multi-task regression bonilla2007multi.
Heterogeneous multiple outputs have also received research focus: while multiple Gausssian-distributed outputs allow for analytical marginalisation of the latent variable, the presence of discrete or other non-Gaussian-distributed outputs forces the use of non-conjugate inference methods such as variational methods, Expectation Propagation or sampling of the latent variable. Here we opt for the latter solution, aided by the development of the Stan programming language and the underlying continuity and smoothness of all latent parameters of interest.
@@ 4 Results
The results describing the behaviour of the trained models are presented here. We ran experiments using 10-fold cross-validation for two GP models with and without the random effect component in the individual-level covariance function ("GP.m" and "GP.f"), and two parametric models run in brms also with or without individual-level random effects ("brms.m" and "brms.f").
@@ 4.1 Comparison in brms
Two brms models were used as comparison methods: ("brms.m" and "brms.f") with or without random effects at the individual level, respectively. Both models had linear fixed effects based on the covariates for each data set, i.e. for the simulated data example, each output modelled with the following model formulae:

### v2

@@ Abstract
Repeated measures, commonly performed in clinical research, facilitate computational acceleration for nonlinear Bayesian nonparametric models and enable exact sampling for non-conjugate inference, when combinations of continuous and discrete endpoints are observed. We make use of Kronecker structure for scaling Gaussian Process models to large-scale, heterogeneous, clinical data sets. Model inference is performed in Stan, and comparisons are made with brms on simulated data and two real clinical data sets, following a radiological image quality theme. Scalable Gaussian Process models compare favourably with parametric models on real data sets with 17,460 observations. Different GP model specifications are explored, with components analogous to random effects, and their theoretical properties are described.
@@ 1 Introduction
Here we present and run new GP models using repeated measurements in the covariate structure and multiple heterogeneous (mixed continuous and discrete) outputs: exact inference in Stanstancore is scaled to tens of thousands of measurements on the processors of a domestic-issue laptop. The models represent the joint covariance between all of the data as Kronecker-structured, and perform Hamiltonian Monte Carlo (HMC) sampling for hyperparameters, missing output values, and on the latent GP space for non-conjugate inference in the presence of heterogeneous outputs. The methods developed here are of use for general application across clinical research, but the real data sets used in this article follow the theme of radiological image quality, in which the covariates consist of patient characteristics, body locations and time, while the outputs represent measurements of image quality, either continuously ”objectively” on the Hounsfield scalehounsfieldunits of radiodensity, or discretely ”subjectively” from expert evaluations.
[…]
@@
  • An implementation of the relevant Gaussian Process models in Stanstancore ; rstan (with wrapper in R), allowing for running the inference on multiple real world data sets, and benchmarking various model specificatons such as covariance function choices.
@@ 2.2 Gaussian Processes
$y|x=$ $f(x)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$
Specific choices of covariance functions $k(x,x^{\prime},\theta)$ have corresponding implicit parametric basis functions. One advantage of the ”function-space” formulation is the ability to use a finite representation of a function with a potentially infinite-dimensional parametric representation: this is the sense in which the models are considered ”nonparametric”, in that they avoid specifying a parametric model for the latent mean function.
@@ 2.3 Scalability and Kronecker-Structure Covariance Matrices
Asserting a GP with a covariance function $k(x,x^{\prime},\theta)$ over a data set with $N$ covariate observations defines a $N\times N$ covariance222also known as kernel matrix matrix $\mathbf{K}$, where the matrix element $\mathbf{K}[i,j]$ is equal to the covariance function $k(x^{i},{x^{j}}^{\prime},\theta)$ evaluated at the $ith$ and $jth$ data point. One challenge of working with GPs is the need to perform a decomposition of the $N\times N$ covariance matrix, which is the dominant computational demand when evaluating the marginal likelihood for sampling or computing the predictive distributions. This results in $\mathcal{O}(N^{3})$ cubic scaling in computational costs with the number of data points $N$, ruling out exact inference for generic covariance matrices for larger data sets . Various methods exist to enable approximate inference for larger data setsliu2020gaussian ; thomas2017scalable , while exact inference is possible for larger data sets when there is structure in the covariance matrix that can be exploited, such as Kronecker or Toeplitzsaatcci2012scalable . In instances where data can be represented as lying in a grid structure, with separable covariance structure between each dimension of the grid, the full covariance matrix can be represented as a Kronecker product between the dimensions. A Kronecker product is an operation performed on two matrices generating a third matrix composed of each individual element of the first matrix separately multiplied with the entire second matrix, combined in a blockwise fashion, i.e. for two matrices $A$ and $B$:
@@
For a data set consisting of covariates measured over a two dimensional grid, for example the pixels of an image, the covariance matrix over the entire data set ($\mathbf{K}_{12}=K([x_{1},x_{2}],[x^{\prime}_{1},x^{\prime}_{2}])$) can be represented as the Kronecker product between two covariance matrices representing each grid dimension independently ($\mathbf{K}_{1}=K(x_{1},x^{\prime}_{1})$, $\mathbf{K}_{2}=K(x_{2},x^{\prime}_{2})$), with the constraint that the covariance functions used are separable, i.e. for a grid composed of $n_{1}$ vertical grid points $x_{1}$, and $n_{2}$ horizontal grid points $x_{2}$, then:
@@
We can apply this reasoning to clinical scenario described in Section 2.1: for a longitudinal study consisting of $N_{1}$ individuals measured at $N_{2}$ time points at $N_{3}$ anatomical locations, the covariate data X can be represented as lying on a three dimensional grid, where one of the grid dimensions $x_{1}$ is the variation between individuals, the second $x_{2}$ as the variation in time, and the third $x_{3}$ variation in anatomical location. Consequently, defining separate covariance matrices representing the variation across individuals (the $N_{1}\times N_{1}$ matrix $\mathbf{K}_{1}$), variation over time (the $N_{2}\times N_{1}$ matrix $\mathbf{K}_{2}$), and anatomical variation (the $N_{3}\times N_{3}$ matrix $\mathbf{K}_{3}$) we can represent the covariance between the covariate measurements $X$ as:
@@ 2.4 Heterogeneous Multi-Task Gaussian Processes
The GP framework extends naturally to multi-dimensional responses $\mathbf{Y}$, analogous to the case of multivariate regressionalvarez2012kernels ; bonilla2007multi . The different response dimensions (also known as ”tasks” or ”outputs”) are appended into a single vector, and a between-task covariance matrix $\mathbf{K}_{out}$ is used to model correlations between the tasks. When the different tasks are evaluated at the same values of the covariates, generating a covariance matrix between covariates $\mathbf{K}_{X}$ then Kronecker structure again emerges in the full covariance matrix $\mathbf{K}_{f}=\mathbf{K}_{out}\otimes\mathbf{K}_{X}$, and computational accelerations become possible.
@@
All of the data sets used in this article use one Kronecker component for the multi-output correlation, and three Kronecker grid components for the covariates, which are grouped into grid dimensions $x_{1}$, $x_{2}$ and $x_{3}$, which may represent variously patient characteristics, intervention treatment, time, body level, or similar. The full covariance structure is therefore:
@@ 2.5 Covariance Matrix Design and Random Effects
Random effects are often desirable in the presence of repeated measurements, where we are not necessarily interested in the variation between individuals but we would still like to include it in the model gelman2007data . This can be performed in an elegant way within GP regression models by including a structured diagonal noise component representing the random variation across individuals. In the Kronecker context, this can be achieved by adding spherical noise to the Kronecker product component that represents the variation between individuals that call for additional random effects, representing the ”noise” sampled when moving between repeated measurements. As this is an additive covariance function where one of the covariance functions is diagonal noise, the same problem emerges when combining additive covariance functions with Kronecker structure: the resulting covariance function components cannot be interpreted totally straightforwardly, as they are multiplied with the covariance functions associated with the other Kronecker components. Formally, a traditional random effects model is represented in the first half of the following inequality, and the model we implemented in the second:
@@
This formulation of a random effect is equivalent to adding some extra variance to the coefficients corresponding to non-patient specific covariates. We detail this in Appendix 7, for the case where all kernels are linear and thus GP regression is equivalent to Bayesian linear regression.
We ran separate models with and without the ”random effect” noise in the individual-level covariance function to explore its influence on the model fit and predictions. We call the models with and without this the mixed-effect GP (”GP.m”) and the fixed-effect GP (”GP.f”).
@@ 2.7 Stan Implementation
The methods described here were implemented in Stanstancore , a probabilistic programming language designed for Bayesian inference, in which model specification is performed explicitly within the language, and Hamiltonian Monte Carlohoffman2014no (HMC) with the No U-Turn Sampler (NUTS) is performed for the model parameters. The Gaussian Process was represented for inference with a standardised, uncorrelated Gaussian latent variable $\eta$, which was reshaped and transformed by the covariance Kronecker components to form the latent variable $f$ conditioned on the covariance. HMC was performed for the GP hyperparameters and noise terms, the latent function representation $\eta$, the latent function $f$ for discrete-valued tasks, and the missing values of the outputs. The Stan programs were called from within the R programming language via Rstanrstan . Similar computational speedups have been achieved in Stan previously flaxman2015fast ; Ronneberg2021 .
@@ 2.8 Missing Data
The inference for continuous-valued missing outputs $y$ can be integrated in a straightforward way with most inference schemes. Given that Stan cannot perform inference over discrete variables, missing binary outputs pose a challenge. This was resolved by calculating the log-likelihood contributions according to the Bernoulli likelihood with an appropriate link function, with the missing $y$ varying continuously between zero and one. This allows for efficient gradient-informed inference procedures, and the interpretation of the imputed variables as probabilities.
@@ 2.9 Radiological Image Quality
While the models described here are of interest across many applications across and beyond clinical research, the application theme of this article is radiological image quality, a field in which it is common to combine continuous-valued ”objective” measurements of image quality with discrete-valued ”subjective” expert evaluations of image quality. This article uses two real data sets, the first of which is observational time-series data, and the second of which comes from a RCT with thorough repeated measures. In each case the effect of primary interest is the effect on the image quality of volume of contrast medium per body weight.
The first real data set was previously published as a clinical study in tran2022indirect . The image quality was measured in venal blood vessels of 53 patients, with continuous-valued attenuation in Hounsfield Units as the objective endpoint, and binary evaluations of image quality from three different consultant radiologists as the subjective endpoints. Each patient was injected with a different quantity of contrast medium per body weight, and evaluations of the images were taken at six thirty-second interval time points after injection. Evaluations were also performed for each side of the body. Further covariates used were gender, age and tube voltage. The combination of 53 patients, 4 endpoints, 7 time points and 2 body sides implies a covariance structure of size 2,968, but the use of the Kronecker product for each of these contributions means that the computations are readily tractable on a personal laptop.
The second real data set comes from a Randomised Control Trial assessing image quality in arterial blood vessels following Computed Tomography Angiography (CTA) kristiansen2023halved . 210 patients were randomised to receive full versus half doses of contrast volume per body weight, with measurements repeated at 7 different locations in the body, and 3 different spectral energy levels measured in kiloelectronVolts (keV). Examples of images recorded at different body levels and energy levels are presented in Figure 1. Attenuation level and image noise recorded in selected Regions Of Interest (ROIs), each measured in Hounsfield units were used as continuous-valued ”objective” outcomes, and two consultant radiologists evaluated every image on a nominal scale for image quality. Because of the very skew distribution of the nominal data, these were simplified to a binary variable of ”Excellent” vs every other level, results into two binary ”subjective” endpoints. Examples of the ordinal scale of subjective image quality are shown in Figure 2. Sex, age and flow speed of injection were also collected as relevant covariates. The combination of 210 patients, 4 endpoints, 7 body locations and three energy levels resulted in a data set of size $N=17,640$ but again the Kronecker structure enabled exact inference with limited computational resources.
@@ 3 Related Work
Gaussian Processes have been used for decades under various names such as kriging or Bayesian kernel regression, and are related to the widely used Support Vector Machines (SVMs) hearst1998support . Their used has increased in the past couple of decades with the advent of greater computational resources, and much research has occurred concerning their scalability to larger data sets under contemporary computational constraints. These methods often include variational inference schemes via the ”inducing point” frameworkhensman2013gaussian , or spectral methods to approximate the full covariance function via sampled Fourier featuresrahimi2007random . The speedups possible through Kronecker structure have been used for some time, but have previously been used on structured image pixel datagilboa2014image , drug combinationsRonneberg2021 ; Ronneberg2023 , spatio-temporal modellingGelfand2004 , and multi-task regression bonilla2007multi .
Heterogeneous multiple outputs have also received research focus moreno2018heterogeneous : while multiple Gaussian-distributed outputs allow for analytical marginalisation of the latent variable, the presence of discrete or other non-Gaussian-distributed outputs forces the use of non-conjugate inference methods such as variational methods hoffman2013stochastic , Expectation Propagation minka2013expectation or sampling of the latent variable posterior. Here we opt for the latter solution, aided by the development of the Stan programming language and the underlying continuity and smoothness of all latent parameters of interest. Heterogeneity between multiple data sets has also been explored with multiple output GPs in the context of fusing randomised and observational data sets dimitriou2024data .
@@ 4 Results
The results describing the behaviour of the trained models are presented here. We ran experiments using 10-fold cross-validation for two GP models with and without the random effect component in the individual-level covariance function (”GP.m” and ”GP.f”), and two parametric models run in brms also with or without individual-level random effects (”brms.m” and ”brms.f”).
@@ 4.1 Comparison in brms
Two brms models were used as comparison methods: (”brms.m” and ”brms.f”) with or without random effects at the individual level, respectively. Both models had linear fixed effects based on the covariates for each data set, i.e. for the simulated data example, each output modelled with the following model formulae: